OpenStudy (anonymous):

Which can not be written as an equivalent improper fraction? fourteen-fourths two and two-thirds one-sixth nine-eighths

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Please Help Anyone I will medal and fan

OpenStudy (anonymous):

https://i.gyazo.com/f783d5cb95bc9de7f1591b9fccd85236.png Hope that helps !

OpenStudy (anonymous):

So your answer would be 1/6

imqwerty (imqwerty):

:)

OpenStudy (jhannybean):

\[\frac{14}{4} ~,~~ \frac{8}{3} ~,~ \frac{1}{6} ~,~ \frac{9}{8}\] An improper fraction is one which the denominator is greater than the numerator, or "top heavy" On the other hand, a proper is where the numerator is smaller than the denominator and its more balanced. Also, an improper fraction can never be written as a proper fraction. So which one IS your proper fraction?

OpenStudy (madhu.mukherjee.946):

1/6

OpenStudy (jhannybean):

A proper fraction can never be written as as improper fraction*